# Transfer of quantum game strategies

### v1

@@ 4. SQNS correlations
4.1. Representations of SQNS correlations via operator systems
Let $\mathcal{S}$ be an operator system. Recall that the universal $C^{*}$-cover of $\mathcal{S}$ (see [27]) is the pair $(C_{u}^{*}(\mathcal{S}),\iota)$ where $C_{u}^{*}(\mathcal{S})$ is a unital ${\rm C}^{*}$-algebra, $\iota:\mathcal{S}\rightarrow C_{u}^{*}(\mathcal{S})$ is a unital complete order embedding such that $\iota(\mathcal{S})$ generates $C_{u}^{*}(\mathcal{S})$ as a ${\rm C}^{*}$-algebra, and whenever $H$ is a Hilbert space with $\phi:\mathcal{S}\rightarrow\mathcal{B}(H)$ a unital completely positive map, there exists a *-representation $\pi_{\phi}:C_{u}^{*}(\mathcal{S})\rightarrow\mathcal{B}(H)$ such that $\pi_{\phi}\circ\iota=\phi$. The following result shows how we can understand the structure of the multivariate commuting tensor product of operator systems when we view them inside the maximal tensor product of their universal ${\rm C}^{*}$-covers.
Theorem 4.10.
Proof.
For the sake of brevity, set $\mathcal{S}={\rm c}-\otimes_{j=1}^{k}\mathcal{S}_{j}$. First, suppose $u\in M_{n}(\mathcal{S})^{+}$ for some $n\in\mathbb{N}$. We wish to show that $u\in M_{n}({\rm max}-\otimes_{j=1}^{k}C_{u}^{*}(\mathcal{S}_{j}))^{+}$. By [26, Lemma 4.1], it is sufficient to prove that $\phi^{(n)}(u)\geq 0$ for each unital completely positive $\phi:{\rm max}-\otimes_{j=1}^{k}C_{u}^{*}(\mathcal{S}_{j})\rightarrow\mathcal{B}(H)$. By Stinespring’s Theorem, we may also without loss of generality assume $\phi$ is a *-homomorphism. By [19, Proposition 7.4], associativity, and the universal property of the maximal tensor product of ${\rm C}^{*}$-algebras, each such $\phi$ is equivalent to $\prod_{j=1}^{k}\pi_{j}$, where $\pi_{j}:C_{u}^{*}(\mathcal{S}_{j})\rightarrow\mathcal{B}(H)$ is a *-homomorphism for $j=1,\ldots,k$ and all have mutually commuting ranges. As the restriction of $\pi_{j}$ to $\mathcal{S}_{j}$ remains completely positive for $j=1,\ldots,k$, the result follows.
Conversely, let $\tau$ be the operator system structure on $\otimes_{j=1}^{k}\mathcal{S}_{j}$ arising from the inclusion $\otimes_{j=1}^{k}\mathcal{S}_{j}\subseteq{\rm max}-\otimes_{j=1}^{k}C_{u}^{*}(\mathcal{S}_{j})$. Suppose that $u\in M_{n}(\tau-\otimes_{j=1}^{k}\mathcal{S}_{j})^{+}$, with $\phi_{j}:\mathcal{S}_{j}\rightarrow\mathcal{B}(H)$ completely positive maps with mutually commuting ranges for $j=1,\ldots,k$. By the definition of $C_{u}^{*}(\mathcal{S}_{j})$, there exists unique *-homomorphisms $\pi_{j}:C_{u}^{*}(\mathcal{S}_{j})\rightarrow\mathcal{B}(H)$ extending $\phi_{j}$, for $j=1,\ldots,k$. As $\mathcal{S}_{j}$ generates $C_{u}^{*}(\mathcal{S}_{j})$ as a ${\rm C}^{*}$-algebra, we may then conclude the ranges of $\pi_{j}$ are mutually commuting as well for $j=1,\ldots,k$. Thus, $(\otimes_{j=1}^{k}\phi_{j})^{(n)}(u)\geq 0$, which implies $u\in M_{n}(\mathcal{S})^{+}$- completing the proof. ∎
Lemma 4.11.
Proof.
We only consider the case when $k=3$; all others will follow similarly. Let $D$ be the set containing all sums of elementary tensors $v_{1}\otimes v_{2}\otimes v_{3}$ with $v_{i}\in V_{i}^{+},\;i=1,2,3$. We claim that if, for every $\epsilon>0$, there exists $u_{\epsilon}\in D$ such that $\|u_{\epsilon}\|\rightarrow 0$ as $\epsilon\rightarrow 0$ and $u+u_{\epsilon}\in D$ for each $\epsilon>0$, then $u\in D$. We may, without loss of generality, further assume that $\|u_{\epsilon}\|\leq 1$ for all $\epsilon>0$. Set $K=2{\rm dim}(V_{1}){\rm dim}(V_{2}){\rm dim}(V_{3})+1$ and, using Carathéodory’s Theorem, start by writing
where $v_{i,j}^{(\epsilon)}\in V_{i}^{+}$ and $\|v_{1,j}^{(\epsilon)}\|=\|v_{2,j}^{(\epsilon)}\|=\|v_{3,j}^{(\epsilon)}\|$ for $i=1,2,3$, $j=1,\ldots,K$ and all $\epsilon>0$. As $v_{1,j}^{(\epsilon)}\otimes v_{2,j}^{(\epsilon)}\otimes v_{3,j}^{(\epsilon)}\leq u+u_{\epsilon}$ and $\|u+u_{\epsilon}\|\leq\|u\|+1$ for all $\epsilon>0$, we get that $\|v_{i,j}^{(\epsilon)}\|\leq\sqrt{\|u\|+1}$ for $i=1,2,3$ and $j=1,\ldots,K$. Using the fact that all of our AOU spaces are finite-dimensional, by compactness we can also assume $v_{i,j}^{(\epsilon)}\rightarrow v_{i,j}$ as $\epsilon\rightarrow 0$ for $i=1,2,3$ and all $j=1,\ldots,K$. We may then conclude that $u=\sum_{j=1}^{K}v_{1,j}\otimes v_{2,j}\otimes v_{3,j}\in D$.
Let
(22) $S_{0}=\sum\limits_{p=1}^{\ell}a_{p}\otimes v_{p}^{(1)},\;\;\;\;T_{0}=\sum\limits_{q=1}^{s}b_{q}\otimes v_{q}^{(2)},\;\;\;\;U_{0}=\sum\limits_{r=1}^{t}c_{r}\otimes v_{r}^{(3)},$
Now suppose first that $S\in M_{n}({\rm OMAX}(V_{1}))^{+}$ and $\alpha\in M_{1,nmd}$. By the description of the multivariate maximal tensor product as provided in [19, Proposition 7.2], if $\epsilon>0$ then $S+\epsilon 1_{n}$ has the form of $S_{0}$ as in (22). Therefore,
$\alpha(S\otimes T_{0}\otimes U_{0})\alpha^{*}\in D.$
For a linear functional
(23) $\big{(}s(e_{x_{2},x_{2}^{\prime},x_{1},x_{1}^{\prime}}\otimes e_{y_{2},y_{2}^{\prime},y_{1},y_{1}^{\prime}}\otimes e_{a_{1},a_{1}^{\prime},a_{2},a_{2}^{\prime}}\otimes e_{b_{1},b_{1}^{\prime},b_{2},b_{2}^{\prime}})\big{)}_{x_{2}x_{2}^{\prime},y_{2}y_{2}^{\prime},a_{1}a_{1}^{\prime},b_{1}b_{1}^{\prime}}^{x_{1}x_{1}^{\prime},y_{1}y_{1}^{\prime},a_{2}a_{2}^{\prime},b_{2}b_{2}^{\prime}}.$
Theorem 4.12.
The map $s\mapsto\Gamma_{s}$ is an affine isomorphism from
  • (i)
  • (ii)
  • (iii)
  • (iv)
Proof.
(i) Let $\Gamma\in\mathcal{Q}_{\rm sns}$. If
for all $x_{2},x_{2}^{\prime}\in X_{2}$. This equality implies
and
Thus,
by the injectivity of the minimal operator system tensor product, $C\in(\mathcal{L}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{L}_{Y_{2},Y_{1}}\otimes_{\rm min}\mathcal{L}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{L}_{B_{1},B_{2}})^{+}$. By [16, Proposition 1.9], and [38, Proposition 5.5],
Arguing now as in [38, Theorem 6.2], this establishes the claim.
(ii) First suppose that
Applying the GNS construction to $\tilde{s}$ and using [38, Theorem 5.2], we obtain a Hilbert space $H$, a unit vector $\xi\in H$ and mutually commuting stochastic operator matrices
satisfying
acting on $K$ so that $\Gamma=\Gamma_{E,F,\eta}$, for $E=E_{X}\cdot E_{Y}$ and $F=F_{A}\cdot F_{B}$. Let $\pi_{X},\pi_{Y},\pi_{A}$ and $\pi_{B}$ be the (unital) *-representations of $\mathcal{C}_{X_{2},X_{1}},\mathcal{C}_{Y_{2},Y_{1}},\mathcal{C}_{A_{1},A_{2}}$ and $\mathcal{C}_{B_{1},B_{2}}$ on $\mathcal{B}(K)$ arising from $E_{X},E_{Y},F_{A}$, and $F_{B}$ respectively (see [38, Theorem 5.2]). Then $\pi:=\pi_{X}\otimes\pi_{Y}\otimes\pi_{A}\otimes\pi_{B}$ is a unital *-representation of $\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm max}\mathcal{C}_{Y_{2},Y_{1}}\otimes_{\rm max}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm max}\mathcal{C}_{B_{1},B_{2}}$ on $\mathcal{B}(K)$. Using Theorem 4.10 once more, if we let $s$ be the restriction to $\mathcal{T}_{X_{2},X_{1}}\otimes_{\rm c}\mathcal{T}_{Y_{2},Y_{1}}\otimes_{\rm c}\mathcal{T}_{A_{1},A_{2}}\otimes_{\rm c}\mathcal{T}_{B_{1},B_{2}}$ of the state
$w\mapsto\langle\pi(w)\eta,\eta\rangle,$
it is clear that $\Gamma=\Gamma_{s}$.
$w\mapsto\langle\pi(w)\tilde{\eta},\tilde{\eta}\rangle,$
Now let $s:\mathcal{T}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{T}_{Y_{2},Y_{1}}\otimes_{\rm min}\mathcal{T}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{T}_{B_{1},B_{2}}\rightarrow\mathbb{C}$ be a state, and using the fact that the minimal operator system tensor product is injective (see [19]) let $\tilde{s}:\mathcal{C}_{X_{2},X_{1}}\otimes_{\rm min}\mathcal{C}_{Y_{2},Y_{1}}\otimes_{\rm min}\mathcal{C}_{A_{1},A_{2}}\otimes_{\rm min}\mathcal{C}_{B_{1},B_{2}}\rightarrow\mathbb{C}$ be an extension of $s$. By [23, Corollary 4.3.10], $\tilde{s}$ can be approximated in the ${\rm weak}^{*}$ topology by elements of the convex hull of vector states on $\pi_{X}(\mathcal{C}_{X_{2},X_{1}})\otimes_{\rm min}\pi_{Y}(\mathcal{C}_{Y_{2},Y_{1}})\otimes_{\rm min}\pi_{A}(\mathcal{C}_{A_{1},A_{2}})\otimes_{\rm min}\pi_{B}(\mathcal{C}_{B_{1},B_{2}})$ (where $\pi_{X},\pi_{Y},\pi_{A},$ and $\pi_{B}$ are unital *-representations of $\mathcal{C}_{X_{2},X_{1}},\mathcal{C}_{Y_{2},Y_{1}},\mathcal{C}_{A_{1},A_{2}}$ and $\mathcal{C}_{B_{1},B_{2}}$, respectively). Using an argument similar to the proof of [7, Theorem 5.6] or [38, Theorem 6.5], we can show that $\Gamma$ is a limit of quantum SQNS correlations.
(iv) This proof is along the lines of the proof for [38, Theorem 6.7]; we include the details for the convenience of the reader. We first let $s$ be a state on
By [25, Theorem 9.9] and [16, Proposition 1.9], we may consider $s$ as an element of
$\bigg{(}{\rm OMAX}(\mathcal{T}_{X_{2},X_{1}})\otimes_{\rm max}{\rm OMAX}(\mathcal{T}_{Y_{2},Y_{1}})\otimes_{\rm max}{\rm OMAX}(\mathcal{T}_{A_{1},A_{2}})\otimes_{\rm max}{\rm OMAX}(\mathcal{T}_{B_{1},B_{2}})\bigg{)}^{+}.$
(24) $\Gamma_{s}=\sum\limits_{j=1}^{k}\lambda_{j}\Phi_{X}^{(j)}\otimes\Phi_{Y}^{(j)}\otimes\Phi_{A}^{(j)}\otimes\Phi_{B}^{(j)}.$
By Remark 4.8, this shows $\Gamma_{s}\in\mathcal{Q}_{\rm sloc}$.
Now suppose $\Gamma$ is of the form (24), with $s$ a functional on $\mathcal{T}_{X_{2},X_{1}}\otimes\mathcal{T}_{Y_{2},Y_{1}}\otimes\mathcal{T}_{A_{1},A_{2}}\otimes\mathcal{T}_{B_{1},B_{2}}$ such that $\Gamma=\Gamma_{s}$. Let $E_{X}^{(j)}\in(M_{X_{2}}\otimes M_{X_{1}})^{+}$ (resp. $E_{Y}^{(j)}\in(M_{Y_{2}}\otimes M_{Y_{1}})^{+},E_{A}^{(j)}\in(M_{A_{1}}\otimes M_{A_{2}})^{+}$, and $E_{B}^{(j)}\in(M_{B_{1}}\otimes M_{B_{2}})^{+}$) be the Choi matrix of $\Phi_{X}^{(j)}$ (resp. $\Phi_{Y}^{(j)},\Phi_{A}^{(j)}$ and $\Phi_{B}^{(j)}$); then $E_{X}^{(j)},E_{Y}^{(j)},E_{A}^{(j)}$, and $E_{B}^{(j)}$ are stochastic operator matrices acting on $\mathbb{C}$. By [38, Theorem 5.2], there exist positive functionals $\phi_{X}^{(j)}:\mathcal{T}_{X_{2},X_{1}}\rightarrow\mathbb{C},\phi_{Y}^{(j)}:\mathcal{T}_{Y_{2},Y_{1}}\rightarrow\mathbb{C},\phi_{A}^{(j)}:\mathcal{T}_{A_{1},A_{2}}\rightarrow\mathbb{C}$, and $\phi_{B}^{(j)}:\mathcal{T}_{B_{1},B_{2}}\rightarrow\mathbb{C}$ such that
and thus, by Lemma 4.11 is a state on
as claimed. ∎
Remark 4.13.
We note that, using an almost identical argument as in (iv), we have an affine isomorphism between the state space of ${\rm OMIN}(\mathcal{S}_{X_{2},X_{1}})\otimes_{\rm min}{\rm OMIN}(\mathcal{S}_{Y_{2},Y_{1}})\otimes_{\rm min}{\rm OMIN}(\mathcal{S}_{A_{1},A_{2}})\otimes_{\rm min}{\rm OMIN}(\mathcal{S}_{B_{1},B_{2}})$ and $\mathcal{C}_{\rm sloc}$; the classical analogue for cases (i)-(iii) are addressed in [19, Theorem 7.11].
Corollary 4.14.
The set $\mathcal{Q}_{\rm sqc}$ is closed and convex.
Proof.
By Theorem 4.12, it is straightforward to show that the affine mapping $s\mapsto\Gamma_{s}$ is also homeomorphism when the state space of $\mathcal{T}_{X_{2}X_{1}}\otimes_{\rm c}\mathcal{T}_{Y_{2}Y_{1}}\otimes_{\rm c}\mathcal{T}_{A_{1}A_{2}}\otimes_{\rm c}\mathcal{T}_{B_{1}B_{2}}$ is equipped with the ${\rm weak}^{*}$-topology. As the state space will be ${\rm weak}^{*}$-compact, the range of this homeomorphism must be (convex and) closed. ∎
Remark 4.15.
By [19, Theorem 7.11], we may use an identical argument as in Corollary 4.14 to show that $\mathcal{C}_{\rm sqc}$ is closed and convex.
Remark 4.16.
Theorem 4.17.
  • (i)
  • (ii)
  • (iii)
  • (iv)
Proof.
Statement (i) follows from Remark 4.16, and an application of [38, Theorem 8.3]. Statement (ii) follows from (i), and Theorem 4.12. Statements (iii) and (iv) follow similarly, when paired with [38, Remark 8.1]. ∎
4.2. SQNS bicorrelations and their representations via operator systems
In the remainder of this subsection, assume that $X_{1}=X_{2}:=X,Y_{1}=Y_{2}:=Y,A_{1}=A_{2}:=A$, and $B_{1}=B_{2}:=B$. Furthermore, assume that $X=A$ and $Y=B$. An SQNS correlation $\Gamma$ will be called an SQNS bicorrelation if $\Gamma$ is unital and $\Gamma^{*}$ is an SQNS correlation. The collection of all SQNS correlations will be denoted $\mathcal{Q}_{\rm sns}^{\rm bi}$. An SQNS bicorrelation $\Gamma$ over $(XY,XY,XY,XY)$ is called quantum commuting if there exists a Hilbert space $H$, strongly bistochastic operator matrices
@@ 4.2. SQNS bicorrelations and their representations via operator systems
with mutually commuting entries acting on $H$, and a unit vector $\xi\in H$ such that $\Gamma=\Gamma_{E,F,\xi}$ where $E=E_{X}\cdot E_{Y},F=F_{A}\cdot F_{B}$; we denote this class by $\mathcal{Q}_{\rm sqc}^{\rm bi}$. The classes of quantum SQNS bicorrelations (denoted $\mathcal{Q}_{\rm sq}^{\rm bi}$), approximately quantum SQNS bicorrelations (denoted $\mathcal{Q}_{\rm sqa}^{\rm bi}$) and local SQNS bicorrelations (denoted $\mathcal{Q}_{\rm sloc}^{\rm bi}$) are defined similarly to their SQNS correlation counterparts, with dilatable strongly bistochastic operator matrices of the appropriate type in place of the strongly stochastic operator matrices of said type. In the sequel, we let $\mathcal{C}_{\rm st}^{\rm bi}$ denote the class of all SQNS bicorrelations of type ${\rm t}$, as defined in [19, Section 6] and Remark 4.5.
@@
As in the setup before Theorem 4.12, for a linear functional
Theorem 4.21.
The map $s\mapsto\Gamma_{s}$ is an affine isomorphism from
  • (i)
  • (ii)
  • (iii)
  • (iv)
Proof.
(i) Let $\Gamma\in\mathcal{Q}_{\rm sns}^{\rm bi}$. If
is the Choi matrix of $\Gamma$ and
for all $x_{i},x_{i}^{\prime}\in X,i=1,2$. Note that
$L_{\rho}(C)\in\mathcal{L}_{X}\text{ for all }\rho\in M_{YAB}.$
$L_{\rho}(C)\in\mathcal{L}_{Y}\text{ for all }\rho\in M_{XAB},$
$L_{\rho}(C)\in\mathcal{L}_{A}\text{ for all }\rho\in M_{XYB},$
and
$L_{\rho}(C)\in\mathcal{L}_{B}\text{ for all }\rho\in M_{XYA}.$
(ii)-(iv) The proofs for (ii)-(iv) follow exactly the same as in the proofs of Theorem 4.12 (ii)-(iv), just by replacing the use of [38, Theorem 5.2] with [7, Theorem 3.4], the use of Remark 4.8 with Remark 4.18, and by replacing $\mathcal{T}_{X_{2},X_{1}}$ and $\mathcal{C}_{X_{2},X_{1}}$ (resp. $\mathcal{T}_{Y_{2},Y_{1}}$ and $\mathcal{C}_{Y_{2},Y_{1}}$, $\mathcal{T}_{A_{1},A_{2}}$ and $\mathcal{C}_{A_{1},A_{2}}$, $\mathcal{T}_{B_{1},B_{2}}$ and $\mathcal{C}_{B_{1},B_{2}}$) with $\mathcal{T}_{X}$ and $\mathcal{C}_{X}$ (resp. $\mathcal{T}_{Y}$ and $\mathcal{C}_{Y}$, $\mathcal{T}_{A}$ and $\mathcal{C}_{A}$, $\mathcal{T}_{B}$ and $\mathcal{C}_{B}$). ∎
We end this subsection with the following (obvious) result.
Corollary 4.22.
The set $\mathcal{Q}_{\rm sqc}^{\rm bi}$ is closed and convex.
@@ 5. Strategy transport
Let $X_{i},Y_{i},A_{i},B_{i}$ be finite sets, for $i=1,2$.
Theorem 5.1.
@@
for $x_{2},x_{2}^{\prime}\in X_{2},a_{2},a_{2}^{\prime}\in A_{2}$. By (2), the former equality implies $L_{\omega_{XA}}(C)\in\mathcal{L}_{Y_{2}B_{2}}$ for every $\omega_{XA}\in M_{X_{2}A_{2}}$, while the latter equality implies $L_{\omega_{YB}}(C)\in\mathcal{L}_{X_{2}A_{2}}$ for every $\omega_{YB}\in M_{Y_{2}B_{2}}$. Thus, $C\in(\mathcal{L}_{X_{2}A_{2}}\otimes\mathcal{L}_{Y_{2}B_{2}})\cap M_{X_{2}Y_{2}A_{2}B_{2}}^{+}$; by the injectivity of the minimal operator system tensor product, $C\in(\mathcal{L}_{X_{2}A_{2}}\otimes_{\rm min}\mathcal{L}_{Y_{2}B_{2}})^{+}$. Argue now as [38, Theorem 6.2] to finish the proof.
@@
(25) $\tilde{E}_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}}=\sum\limits_{x_{1},x_{1}^{\prime}\in X_{1}}\sum\limits_{a_{1},a_{1}^{\prime}\in A_{1}}P_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}Q_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}\otimes E_{x_{1}x_{1}^{\prime},a_{1}a_{1}^{\prime}}$
@@
(26) $\tilde{F}_{y_{2}y_{2}^{\prime},b_{2}b_{2}^{\prime}}=\sum\limits_{y_{1},y_{1}^{\prime}\in Y_{1}}\sum\limits_{b_{1},b_{1}^{\prime}\in B_{1}}P^{y_{2}y_{2}^{\prime},y_{1}y_{1}^{\prime}}Q^{b_{1}b_{1}^{\prime},b_{2}b_{2}^{\prime}}\otimes F^{y_{1}y_{1}^{\prime},b_{1}b_{1}^{\prime}}.$
@@
${\rm Tr}_{A_{2}}(\tilde{E})$ $=$ $\sum\limits_{a_{2}}\sum\limits_{x_{2},x_{2}^{\prime}}\sum\limits_{x_{1},x_{1}^{\prime}}\sum\limits_{a_{1},a_{1}^{\prime}}P_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}Q_{a_{1}a_{1}^{\prime},a_{2}a_{2}}\otimes E_{x_{1}x_{1}^{\prime},a_{1}a_{1}^{\prime}}$
@@
6. Perfect strategies for quantum games
For an arbitrary Hilbert space $H$, we write $\overline{H}$ for the Banach space dual of $H$; by the Riesz Representation Theorem, there exists a conjugate linear isometry $\partial:H\rightarrow\overline{H}$, such that $\partial(\xi)(\eta)=\langle\eta,\xi\rangle,\;\xi,\eta\in H$. In what follows, we will write ¯$\overline{\mathbb{C}}^{X}=\overline{\mathbb{C}^{X}}$ for any finite set $X$. Set $\overline{\xi}=\partial(\xi)$, for $\xi\in H$. Given a linear operator $A:H\rightarrow K$, (where $H,K$ are Hilbert spaces) let $\overline{A}:\overline{K}\rightarrow\overline{H}$ be its (Banach space) dual operator. Note that $\overline{A}(\overline{\xi})=\overline{A^{*}\xi}$, for $\xi\in K$.

### v2

@@ A $\it classical\;strongly\;no-signalling\;(SNS)\;correlation$} [20] is a correlation
In the remainder of this section, assume that $X_{1}=X_{2}:=X,Y_{1}=Y_{2}:=Y,A_{1}=A_{2}:=A$, and $B_{1}=B_{2}:=B$. Furthermore, assume that $X=A$ and $Y=B$. An SQNS correlation $\Gamma$ will be called an SQNS bicorrelation if $\Gamma$ is unital and $\Gamma^{*}$ is an SQNS correlation. The collection of all SQNS correlations will be denoted $\mathcal{Q}_{\rm sns}^{\rm bi}$. An SQNS bicorrelation $\Gamma$ over $(XY,XY,XY,XY)$ is called quantum commuting if there exists a Hilbert space $H$, strongly bistochastic operator matrices
@@
with mutually commuting entries acting on $H$, and a unit vector $\xi\in H$ such that $\Gamma=\Gamma_{E,F,\xi}$ where $E=E_{X}\cdot E_{Y},F=F_{A}\cdot F_{B}$; we denote this class by $\mathcal{Q}_{\rm sqc}^{\rm bi}$. The classes of quantum SQNS bicorrelations (denoted $\mathcal{Q}_{\rm sq}^{\rm bi}$), approximately quantum SQNS bicorrelations (denoted $\mathcal{Q}_{\rm sqa}^{\rm bi}$) and local SQNS bicorrelations (denoted $\mathcal{Q}_{\rm sloc}^{\rm bi}$) are defined similarly to their SQNS correlation counterparts, with dilatable strongly bistochastic operator matrices of the appropriate type in place of the strongly stochastic operator matrices of said type. In the sequel, we let $\mathcal{C}_{\rm st}^{\rm bi}$ denote the class of all SQNS bicorrelations of type $\rm t$}, as defined in [20, Section 6] and Remark 3.8.
@@ 4. Strategy transport
Let $X_{i},Y_{i},A_{i},B_{i}$ be finite sets, for $i=1,2$. Recall that if $\mathcal{E}$ is a QNS correlation over $(X_{1},Y_{1},A_{1},B_{1})$, it acts as a quantum channel $\mathcal{E}:M_{X_{1}Y_{1}}\rightarrow M_{A_{1}B_{1}}$; thus, if $\Gamma$ is SQNS over $(X_{2}Y_{2},A_{1}B_{1},X_{1}Y_{1},A_{2}B_{2})$, then $\Gamma[\mathcal{E}]:M_{X_{2}Y_{2}}\rightarrow M_{A_{2}B_{2}}$. The two theorems in this section are crucial to our goal of transferring strategies between quantum games; briefly, they say that if $\mathcal{E}$ is viewed as a strategy of some type for one game, under mild conditions on $\Gamma$ the simulated channel $\Gamma[\mathcal{E}]$ can be viewed as a strategy of the same type for another quantum game.
Theorem 4.1
@@
for $x_{2},x_{2}^{\prime}\in X_{2},a_{2},a_{2}^{\prime}\in A_{2}$. By (8), the former equality implies $L_{\omega_{XA}}(C)\in\mathcal{L}_{Y_{2}B_{2}}$ for every $\omega_{XA}\in M_{X_{2}A_{2}}$, while the latter equality implies $L_{\omega_{YB}}(C)\in\mathcal{L}_{X_{2}A_{2}}$ for every $\omega_{YB}\in M_{Y_{2}B_{2}}$. Thus, $C\in(\mathcal{L}_{X_{2}A_{2}}\otimes\mathcal{L}_{Y_{2}B_{2}})\cap M_{X_{2}Y_{2}A_{2}B_{2}}^{+}$; by the injectivity of the minimal operator system tensor product, $C\in(\mathcal{L}_{X_{2}A_{2}}\otimes_{\rm min}\mathcal{L}_{Y_{2}B_{2}})^{+}$. Argue now as [39, Theorem 6.2] to finish the proof.
@@
$\tilde{E}_{x_{2}x_{2}^{\prime},a_{2}a_{2}^{\prime}}=\sum\limits_{x_{1},x_{1}^{\prime}\in X_{1}}\sum\limits_{a_{1},a_{1}^{\prime}\in A_{1}}P_{x_{2}x_{2}^{\prime},x_{1}x_{1}^{\prime}}Q_{a_{1}a_{1}^{\prime},a_{2}a_{2}^{\prime}}\otimes E_{x_{1}x_{1}^{\prime},a_{1}a_{1}^{\prime}}$ (22)
@@
$\tilde{F}_{y_{2}y_{2}^{\prime},b_{2}b_{2}^{\prime}}=\sum\limits_{y_{1},y_{1}^{\prime}\in Y_{1}}\sum\limits_{b_{1},b_{1}^{\prime}\in B_{1}}P^{y_{2}y_{2}^{\prime},y_{1}y_{1}^{\prime}}Q^{b_{1}b_{1}^{\prime},b_{2}b_{2}^{\prime}}\otimes F^{y_{1}y_{1}^{\prime},b_{1}b_{1}^{\prime}}.$ (23)
@@
[…]
@@
5. Perfect strategies for various quantum games
For an arbitrary Hilbert space $H$, we write $\overline{H}$ for the Banach space dual of $H$; by the Riesz Representation Theorem, there exists a conjugate linear isometry $\partial:H\rightarrow\overline{H}$, such that $\partial(\xi)(\eta)=\langle\eta,\xi\rangle,\;\xi,\eta\in H$. In what follows, we will write $\overline{\mathbb{C}}^{X}=\overline{\mathbb{C}^{X}}$ for any finite set $X$. Set $\overline{\xi}=\partial(\xi)$, for $\xi\in H$. Given a linear operator $A:H\rightarrow K$, (where $H,K$ are Hilbert spaces) let $\overline{A}:\overline{K}\rightarrow\overline{H}$ be its (Banach space) dual operator. Note that $\overline{A}(\overline{\xi})=\overline{A^{*}\xi}$, for $\xi\in K$.